# Non-relativistic transport from frame-indifferent kinetic theory

### v1

@@ IV Kinetic theory and conservation equations
Kinetic theory models the statistical behavior of large systems of particles, typically in gases, by studying their microscopic motions. The central object in kinetic theory is the phase space distribution function $f(x,p)$, which gives the number of particles $dN$ contained in a space-like volume at point $x$ and having momenta within a small range around $p$. Since the number of particles is an observable, the phase space volume needs to be invariant under all the coordinate and frame ambiguities of NC spaces. In other words, $dN=f\,\Omega_{2d}$ for some invariant volume $2d$-form $\Omega_{2d}$, which we will now construct. To begin with, notice that the bi-linear form $G_{\mu\nu}$ induces a volume form on the phase space
@@
is a volume element on the tangent space to the point $x$ and $\omega_{d+1}$ is the spacetime volume formed introduced in Eq. (9). This is analogous to the relativistic case [66, 67, 68]. Another important ingredient in kinetic theory is Liouville’s theorem, which guarantees the conservation of the flux of trajectories crossing bounded space-like domains of the manifold. Therefore, it is necessary that $\Omega_{2d}$ satisfies £$\pounds_{X_{L}}\Omega_{2d}=0$ and $d\Omega_{2d}=0$. In fact, the following form satisfies such properties [69]
@@
IV.1 Thermal NC spacetime, detailed balance and thermal equilibrium
In [57] NC spacetimes were generalized by adding to the manifold an extra timelike vector field $u^{\mu}$, which without loss of generality can be normalized as $u^{\mu}\tau_{\mu}=1$. As we will see later, such a field can always be defined in the presence of a gas, in which case $u^{\mu}$ describes the velocity of the gas. For such spaces we could argue that the “Milne symmetry” is spontaneously broken since $u^{\mu}$ introduces a distinguished Milne frame with velocity $v_{u}^{\mu}=u^{\mu}$. Notice that this frame is connected to an arbitrary one via a Milne boost $v^{\mu}-v_{u}^{\mu}=h^{\mu\nu}\psi_{\nu}$. The presence of $u^{\mu}$ allows us to define the invariant spatial metric $g_{\mu\nu}$ and $U(1)$ gauge field $A_{\mu}$ as
$A_{\mu}$ $=m_{\mu}+u_{\mu}-\frac{u^{2}}{2}\tau_{\mu}\,,$ (40)
where $u_{\mu}=h_{\mu\nu}u^{\nu}$ and $u^{2}=h_{\mu\nu}u^{\mu}u^{\nu}$. The objects $g_{\mu\nu}$ and $A_{\mu}$ are precisely equal to $h_{\mu\nu}$ and $m_{\mu}$ evaluated in the frame $v_{u}^{\mu}$. We introduce the invariant spatial projector and an affine connection corresponding to $v_{u}^{\mu}$ as, respectively, [57]
with
$f=dA\,.$ (43)
A comparison with Eq. (7) reveals that $\bar{\Gamma}^{\mu}_{\alpha\beta}=\Gamma^{\mu}_{\alpha\beta}$ in the torsionless case. The covariant derivative of the velocity field can be decomposed as
¯$\bar{\nabla}_{\mu}u^{\nu}=\tau_{\mu}a^{\nu}+\frac{1}{2}g_{\mu\alpha}\omega^{\alpha\nu}+g_{\mu\alpha}\sigma^{\alpha\nu}+\frac{1}{d}\bar{P}_{\mu}^{\nu}\theta\,,$ (44)
$a^{\mu}$ $\equiv u^{\nu}\bar{\nabla}_{\nu}u^{\mu}=-f^{\mu}\,_{\nu}u^{\nu}\,,$ (45)
$\sigma^{\mu\nu}$ $\equiv P^{\mu\nu\alpha}\,_{\lambda}\bar{\nabla}_{\alpha}u^{\lambda}\,,$ (47)
$\theta$$\equiv\bar{\nabla}_{\mu}u^{\mu}\,,$ (48)
where
$P^{\mu\nu\alpha}\,_{\lambda}=\frac{1}{2}\left(\delta^{\mu}_{\lambda}h^{\nu\alpha}+\delta^{\nu}_{\lambda}h^{\mu\alpha}-\frac{2}{d}h^{\mu\nu}\delta^{\alpha}_{\lambda}\right)\,.$ (49)
The new structure added to the NC manifold may seem unneccessary. However, in what follow we will prove that the existence of the notion of thermal equilibrium is only possible on this class of NC spaces. Therefore, we will call them Thermal Newton-Cartan (TNC) manifolds.
$S=-\int_{V\times T_{V}}f\log f\,\Omega_{2d}\,,$ (50)
$\dot{S}=X_{L}[S]\geq 0\,.$ (51)
We define a local equilibrium state by a distribution function $f_{0}$ that saturates the inequality in Eq. (51), i.e., $\dot{S}_{0}=0$. This happens if $\log f_{0}$ is a collisional invariant, so that $f_{0}$ obeys the detailed balance condition $\mathcal{C}[f_{0}]=0$. We will call $f_{0}$ a hydrodynamic distribution function. According to the above discussion,
$f_{0}(x,p)\equiv\mathcal{N}\exp\left[Q_{0}(x,p)\right]\,,$ (52)
$Q_{0}=\alpha(x)+\xi^{\mu}(x)h_{\mu\nu}p^{\nu}-\gamma(x)p^{2}\,,$ (53)
where $\alpha(x)$, $\gamma(x)$, and $\xi^{\mu}(x)$ are arbitrary functions constrained by the requirement that $Q_{0}$ must be a scalar. By completing the square $Q_{0}$ can be expressed as
$Q_{0}=-\frac{1}{2mT}(p-mu)^{2}+\frac{\mu}{T}\,,$ (54)
[…]
$T=\frac{1}{\tau_{\mu}\xi^{\mu}}\,,~{}~{}~{}~{}u^{\mu}=T\xi^{\mu}\,.$ (55)
Finally, the collisional invariant can be expressed in the form of Eq. (33)
$Q_{0}=\xi^{\mu}\pi_{\mu}+m\Lambda\,,~{}$ (56)
with the gauge parameter $\Lambda$ satisfying the relation
$\frac{\mu}{mT}=\xi^{\mu}A_{\mu}+\Lambda\,.$ (57)
$X_{L}[f_{\mathrm{eq}}]=0\,.$ (58)
Following an explicit calculation of $X_{L}[f_{0}]$ in Appendix B.2 we find that
which is equivalent to requiring that $\chi=\chi_{K}$ with $\chi_{K}$ being a Killing transformation, see Eq. (30). Thus, $f_{\mathrm{eq}}=\exp\left[Q_{K}(x,p)\right]$. In fact, the Killing conditions in Eq. (60) can be equivalently expressed in terms of the local thermodynamic variables $(T,u^{\mu},\mu)$ as hydrostatic conditions derived in Appendix B.3:
$\theta=0\,,\quad\quad\qquad\sigma^{\mu\nu}=0\,.$ (63)
IV.2 Currents and conservation laws
@@
with $T_{x}$ the tangent space at the point $x$. Associated to the observer with velocity $v^{\mu}$ we can define the amount of kinematic momentum $P_{\mu}$ and energy contained in a volume $V$
$P_{\mu}$ $=\int_{V\times T_{V}}h_{\mu\nu}p^{\nu}f\,\Omega_{2d}=\int_{V}h_{\mu\alpha}\mathcal{T}^{\alpha\nu}(x)\Sigma_{\nu}\,,$ (66)
$E$ $=\frac{1}{2m}\int_{V\times T_{V}}h_{\mu\nu}p^{\mu}p^{\nu}\,f\,\Omega_{2d}=\int_{V}\left(u_{\alpha}-\frac{1}{2}u^{2}\tau_{\alpha}\right)\mathcal{T}^{\mu\alpha}\Sigma_{\mu}+\int_{V}\mathcal{E}^{\mu}(x)\Sigma_{\mu}\,,$ (67)
with the corresponding currents defined as
@@
Notice that particle current $\mathcal{J}^{\mu}$ and momentum density $\mathcal{P}_{\mu}=h_{\mu\alpha}\mathcal{\tau}_{\nu}\mathcal{T}^{\alpha\nu}$ obey the Milne Ward identity
$\mathcal{P}_{\mu}=mh_{\mu\nu}\mathcal{J}^{\nu}\,,$ (70)
and the energy current $\mathcal{E}^{\mu}$ has been defined such that the observer-ambiguous contribution has been subtracted. The energy current $\mathcal{E}^{\mu}$ is Milne-invariant and it can be interpreted as the energy flux in the fluid’s comoving frame.
Typically in kinetic theory it is useful to introduce the higher moments of the distribution function, which in our NC description can be defined as
@@
(¯$\begin{split}\left(\bar{\nabla}_{\mu}+H_{\mu\nu}u^{\nu}\right)\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=&\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}X_{L}[f]\sigma_{d+1}\delta(\tau_{\mu}p^{\mu}-m)\\ &-\frac{n-1}{2}H^{(\alpha_{2}}_{~{}~{}~{}\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\alpha_{3}\cdots\alpha_{n})}\,.\end{split}$ (72)
@@
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\nu}u^{\nu}\right)\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\mathcal{K}^{\alpha_{2}\cdots\alpha_{n}}-\frac{n-1}{2}H^{(\alpha_{2}}\,_{\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\alpha_{3}\cdots\alpha_{n})}\,,$ (73)
@@
Thus, the equations of motion for the mass, momentum and energy currents reduce to the set of conservation equations (the derivation of the equation for $\mathcal{E}^{\mu}$ can be found in Appendix B.5):
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{J}^{\mu}$ $=0\,,$ (77)
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{T}^{\mu\nu}$ $=-H^{\nu}_{~{}~{}\rho}\mathcal{E}^{\rho}\,,$ (78)
(¯$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}$ $=-H_{\mu\rho}u^{\rho}\mathcal{E}^{\mu}-\mathcal{T}^{\mu\nu}g_{\rho(\mu}\bar{\nabla}_{\nu)}u^{\rho}\,.$ (79)
Remarkably, the momentum and energy conservation equations (78), (79) show energy-related “Lorentz force” and “Joule heating” respectively, consistent with the single particle equations of motion (23). Moreover, Eqs. (77), (78), and (79) agree with the Ward identities derived in Refs. [38, 57] using a field-theoretical approach, showing the reliability of our kinetic theory.

### v2

@@ IV Kinetic theory and conservation equations
Kinetic theory models the statistical behavior of large systems of particles, typically in gases, by studying their microscopic motion. The central object in kinetic theory is the phase space distribution function $f(x,p)$ , which gives the number of particles $dN$ contained in a space-like volume at point $x$ and having momenta within a small range around $p$. Since the number of particles is an observable, the phase space volume needs to be invariant under all the coordinate and frame ambiguities of NC spaces. In other words, $dN=f\,\Omega_{2d}$ for some invariant volume $2d$-form $\Omega_{2d}$, which we will now construct. To begin with, notice that the bi-linear form $G_{\mu\nu}$ induces a volume form on the phase space
@@
is a volume element on the tangent space to the point $x$ and $\omega_{d+1}$ is the spacetime volume formed introduced in Eq. (10). This is analogous to the relativistic case [68, 69, 70]. Another important ingredient in kinetic theory is Liouville’s theorem, which guarantees the conservation of the flux of trajectories crossing bounded space-like domains of the manifold. Therefore, it is necessary that $\Omega_{2d}$ satisfies $\pounds_{X_{L}}\Omega_{2d}=0$ and $d\Omega_{2d}=0$. In fact, the following form satisfies such properties [71]
@@
IV.1 Currents and conservation laws
@@
with $T_{x}$ the tangent space at the point $x$. Associated to the observer with velocity $v^{\mu}$, we can define the amount of kinematic momentum $P_{\mu}$ and energy contained in a volume $V$
$P_{\mu}$ $=\int_{V\times T_{V}}p_{\mu}f\,\Omega_{2d}=\int_{V}\mathcal{(}\mathcal{P}_{\mu}v^{\nu}+T_{\mu\alpha}h^{\alpha\nu})\Sigma_{\nu}\equiv\int_{V}\mathcal{T}_{\mu}\,^{\nu}(x)\Sigma_{\nu}\,$, (56)
$E$ $=\frac{1}{2m}\int_{V\times T_{V}}p^{2}\,f\,\Omega_{2d}=\int_{V}\tilde{\mathcal{E}}^{\mu}(x)\Sigma_{\mu}\equiv\int_{V}\mathcal{E}^{\mu}(x)\Sigma_{\mu}+\int_{V}\left(u_{\alpha}-\frac{1}{2}u^{2}\tau_{\alpha}\right)\mathcal{T}^{\mu\alpha}\Sigma_{\mu}\,$, (57)
where the corresponding currents are defined as
@@
When writing Eqs. (57) and (59), we exploited the fact that for matter at non-zero density there always exists a velocity field $u^{\mu}$ associated to the rest frame, so that the spacetime has the structure of a TNC manifold described in Sec. II.2. Consequently, the energy current $\mathcal{E}^{\mu}$ has been defined so that the observer-ambiguous contribution has been subtracted. The energy current $\mathcal{E}^{\mu}$ is Milne invariant and can be interpreted as the energy flux in the fluid’s co-moving frame. Moreover, we have decomposed the invariant “thermal” stress-energy tensor $\mathcal{T}^{\mu\nu}$ into the frame-dependent momentum density $\mathcal{P}_{\mu}$ and the stress tensor $T_{\mu\nu}$. The invariant and frame-dependent currents are related as follows:
$\mathcal{T}^{\mu\nu}$ $=h^{\mu\alpha}h^{\nu\beta}T_{\alpha\beta}+h^{\mu\alpha}\mathcal{P}_{\alpha}v^{\nu}+h^{\nu\alpha}\mathcal{P}_{\alpha}v^{\mu}+m(\tau_{\rho}\mathcal{J}^{\rho})\,v^{\mu}v^{\nu}\,$, (60)
$\tilde{\mathcal{E}}^{\mu}$ $=\mathcal{E}^{\mu}+\mathcal{T}^{\mu\alpha}\left(u_{\alpha}-\frac{1}{2}u^{2}\tau_{\alpha}\right)\,.$ (61)
As we show below, in the hydrodynamic regime, all the microscopic information of the system is encoded in constitutive relations for the invariant currents. Also notice that the particle current and momentum density obey the Milne Ward identity
$\mathcal{P}_{\mu}=mh_{\mu\nu}\mathcal{J}^{\nu}\,.$ (62)
Typically, in kinetic theory it is useful to introduce the higher moments of the distribution function, which in our NC description can be defined as
@@
$\begin{split}\left(\bar{\nabla}_{\mu}+H_{\mu\nu}u^{\nu}\right)\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=&\frac{1}{m^{n-1}}\int p^{\alpha_{2}}\cdots p^{\alpha_{n}}X_{L}[f]\sigma_{d+1}\delta(\tau_{\mu}p^{\mu}-m)\\ &-\frac{n-1}{2}H^{(\alpha_{2}}_{~~~\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\alpha_{3}\cdots\alpha_{n})}\,.\end{split}$ (64)
@@
$\left(\bar{\nabla}_{\mu}+H_{\mu\nu}u^{\nu}\right)\mathcal{I}^{\mu\alpha_{2}\cdots\alpha_{n}}=\mathcal{K}^{\alpha_{2}\cdots\alpha_{n}}-\frac{n-1}{2}H^{(\alpha_{2}}\,_{\rho}g_{\beta\gamma}\mathcal{I}^{\beta\gamma\rho|\alpha_{3}\cdots\alpha_{n})}\,$, (65)
@@
Thus, the equations of motion for the mass, momentum and energy currents reduce to the set of conservation equations (with the derivation of the equation for $\mathcal{E}^{\mu}$ relegated to Appendix B.5):
$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{J}^{\mu}$ $=0\,$, (69)
$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{T}^{\mu\nu}$ $=-H^{\nu}_{~~\rho}\mathcal{E}^{\rho}\,$, (70)
$\left(\bar{\nabla}_{\mu}+H_{\mu\rho}u^{\rho}\right)\mathcal{E}^{\mu}$ $=-H_{\mu\rho}u^{\rho}\mathcal{E}^{\mu}-\mathcal{T}^{\mu\nu}g_{\rho(\mu}\bar{\nabla}_{\nu)}u^{\rho}\,.$ (71)
We see that (69) and (70) are not independent since the former is obtained after contracting the latter with the clock form $\tau_{\mu}$. In addition, the momentum and energy conservation equations (70), (71), show energy-related “Lorentz force” and “Joule heating”, respectively, consistent with the single-particle equations of motion (38). Most importantly, Eqs. (69), (70), and (71) agree with the Ward identities derived in Refs. [38, 58] using a field-theoretical approach, showing the reliability of our kinetic theory. We emphasize Eqs. (69), (70) and (71) contain the same information as the set of equations that we could have obtained if we had worked with the frame-dependent currents.
IV.2 Detailed balance and thermal equilibrium
$S=-\int_{V\times T_{V}}f\log f\,\Omega_{2d}\,$, (72)
$\dot{S}=X_{L}[S]\geq 0\,.$ (73)
We define a local equilibrium state by a distribution function $f_{0}$ that saturates the inequality in Eq. (73), i.e., $\dot{S}=0$. This happens if $\log f_{0}$ is a collisional invariant, so that $f_{0}$ obeys the detailed balance condition $\mathcal{C}[f_{0}]=0$. We will call $f_{0}$ a hydrodynamic distribution function. According to the above discussion,
$f_{0}(x,p)\equiv\mathcal{N}\exp\left[Q_{0}(x,p)\right]\,$, (74)
$Q_{0}=\alpha(x)+\xi^{\mu}(x)h_{\mu\nu}p^{\nu}-\gamma(x)p^{2}\,$, (75)
where $\alpha(x)$, $\gamma(x)$, and $\xi^{\mu}(x)$ are arbitrary functions constrained by the requirement that $Q_{0}$ must be a scalar. By completing the square, $Q_{0}$ can be expressed as
$Q_{0}=-\frac{1}{2mT}(p-mu)^{2}+\frac{\mu}{T}\,$, (76)
where $2mT=\gamma^{-1}$, $u^{\mu}=T\xi^{\mu}$, and $\mu=T\alpha+mu^{2}/2$. The invariance of $Q_{0}$ under Milne boost fixes $u^{\mu}\tau_{\mu}=1$ [^6] Under Milne boost $\delta Q_{0}\propto(p^{\mu}-mu^{\mu})\tau_{\mu}$, which has to vanish on the mass shell., which implies the useful relations
$T=\frac{1}{\tau_{\mu}\xi^{\mu}}\,,~~~~u^{\mu}=T\xi^{\mu}\,.$ (77)
Finally, the collisional invariant can be expressed in the form of Eq. (48),
$Q_{0}=\xi^{\mu}\pi_{\mu}+m\Lambda\,$,~ (78)
with the gauge parameter $\Lambda$ satisfying the relation
$\frac{\mu}{mT}=\xi^{\mu}A_{\mu}+\Lambda\,.$ (79)
We interpret the parameters $T,u^{\mu},\mu$ as hydrodynamic temperature, fluid velocity and chemical potential, respectively.
$X_{L}[f_{\mathrm{eq}}]=0\,.$ (80)
Following an explicit calculation of $X_{L}[f_{0}]$ in Appendix B.2, we find that
which is equivalent to requiring that $\chi=\chi_{K}$ with $\chi_{K}$ being a Killing transformation, see Eq. (45). Thus, $f_{\mathrm{eq}}=\exp\left[Q_{K}(x,p)\right]$. In fact, the Killing conditions in Eq. (82) can be equivalently expressed in terms of the local thermodynamic variables $(T,u^{\mu},\mu)$ as hydrostatic conditions, derived in Appendix B.4:
$\theta=0\,,\quad\quad\qquad\sigma^{\mu\nu}=0\,.$ (85)
Note that temperature, fluid velocity, and chemical potential are well defined variables only at equilibrium where they are entirely fixed by the isometries of the space-time, and that on manifolds without time-like Killing symmetry equilibration is not possible.